OpenStudy (anonymous):

1. Show that absolute value of r

OpenStudy (queelius):

Part 1: abs(r) = { if r < 0, then -r, otherwise r } r < s => if r < 0, then -r < s, multiply by the factor, -1, on both sides => r > -s if r > 0, then r < s. Combining these two inequalities, we see that -s < r < s. Part 2: We are given that s is a non-zero positive rational number. r > -s is true for any positive s if r is zero or any positive rational. If r < 0, then it is possible to choose an r smaller than -s. So, let us only consider zero or positive rational numbers for r. r < s is true for any positive s if r is zero, otherwise you can pick an s smaller than r, e.g., let r = q/p. Then, choose s = (q-1)/p. Thus, the only r which satisfies the inequality for any s is r = 0.